Starting from rest, a 1.9x10-4 kg flea springs straight upward. While the flea is pushing off from the ground, the ground exerts an average upward force of 0.36 N on it. This force does 2.0x10-4 J of work on the flea. (a) What is the flea's speed when it leaves the ground? (b) How far upward does the flea move while it is pushing off? Ignore both air resistance and the flea's weight.
Answer #1

Mass of the flea = m = 1.9 x 10-4 kg

Average upward force on the flea = F = 0.36 N

Work done by the force on the flea = W = 2 x 10-4 J

Speed of the flea when it leaves the ground = V

The work done by the force on the flea is equal to the kinetic energy gained by the flea.

W = mV2/2

2x10-4 = (1.9x10-4)V2/2

V = 1.45 m/s

Distance the flea moves upwards while pushing off = D

The work done by a force is equal to the force times the displacement.

W = FD

2x10-4 = (0.36)D

D = 5.555 x 10-4 m

a) Speed of the flea when it leaves the ground = 1.45 m/s

b) Distance the flea moves upwards while pushing off = 5.555 x 10-4 m